OpenStudy (anonymous):

if 20 percent of 100 is equal to 500 percent of n, then n is equal to what number?

OpenStudy (anonymous):

20% of 100 = (20/100)*100 = 0.2*100 = 20 500% of 4 = (500/100)*4 = 5*4 = 20

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

@abayomi12

OpenStudy (anonymous):

0.2*100=20= 5*n n=4

OpenStudy (anonymous):

(20/100)*100 = 500n/100 20 = 5n n = 4

OpenStudy (anonymous):

ohhhhhhh okk thanks you to